OpenStudy (anonymous):

Find a quadratic model for the set of values. x -2, 0,4 f(x) 1, -3 , 85

OpenStudy (anonymous):

Okay well we know it is going to be in the form:\[f(x)=ax^2+bx+c\]Right? Now each point gives us a new equation:\[1=a(-2)^2+b(-2)+c\]\[-3=a(0)^2+b(0)+c\]\[85=a(4)^2+b(4)+c\]So we have a system of equations... 3 equation and 3 unknowns. Can you solve for \(a\), \(b\), and \(c\)? Do you need help solving this system?

OpenStudy (anonymous):

help please! :(

OpenStudy (anonymous):

Okay.... So the second equation looks most easy to start with so:\[-3=a(0)^2+b(0)+c\]\[0 -3=+ 0 + c\]\[-3=c\]Now we have \(c\).

OpenStudy (anonymous):

We can now make our system a bit simpler:\[1=a(-2)^2+b(-2)-3\]\[85=a(4)^2+b(4)-3\]

OpenStudy (anonymous):

Do you understand what I've done so far?

OpenStudy (anonymous):

somewhat I do..

OpenStudy (anonymous):

What course are you taking right now?

OpenStudy (anonymous):

algebra2

OpenStudy (anonymous):

Have you ever had to solve a system of equations... like: \[x+y=4\]\[4x-y=1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well what I gave you earlier is just a system of equations, but the variables are a, b, and c!

OpenStudy (anonymous):

You use the same methods to solve for a, b, and c.

OpenStudy (anonymous):

wait im really confused

OpenStudy (anonymous):

okay.... let's just pretend that we're using x, y and z to make things simpler...

OpenStudy (anonymous):

\[1=x(-2)^2+y(-2)+z\]\[-3=x(0)^2+y(0)+z\]\[85=x(4)^2+y(4)+z\] So first let's get rid of the exponents and parenthesis.\[1=4x-2y+z\]\[-3=z\]\[85=16x+4y+z\]Can you solve this system of equations?